# MiniConGTS: A Near Ultimate Minimalist Contrastive Grid Tagging Scheme for Aspect Sentiment Triplet Extraction

### v1

@@ Abstract
Aspect Sentiment Triplet Extraction (ASTE) aims to co-extract the sentiment triplets in a given corpus. Existing approaches within the pretraining-finetuning paradigm tend to either meticulously craft complex tagging schemes and classification heads, or incorporate external semantic augmentation to enhance performance. In this study, we, for the first time, re-evaluate the redundancy in tagging schemes and the internal enhancement in pretrained representations. We propose a method to improve and utilize pretrained representations by integrating a minimalist tagging scheme and a novel token-level contrastive learning strategy. The proposed approach demonstrates comparable or superior performance compared to state-of-the-art techniques while featuring a more compact design and reduced computational overhead. Additionally, we are the first to formally evaluate GPT-4’s performance in few-shot learning and Chain-of-Thought scenarios for this task. The results demonstrate that the pretraining-finetuning paradigm remains highly effective even in the era of large language models.
@@ 1 Introduction
Aspect-Based Sentiment Analysis (ABSA) aims to jointly extract opinion terms, aspect terms (targets of the corresponding opinions), and their specific sentiment polarities in a given corpus. In the milestone research by Peng et al. (2020), the compound ABSA subtasks were consolidated into the Aspect Sentiment Triplet Extraction (ASTE) framework. For each input corpus, ASTE outputs triplets in the form (Aspect, Opinion, Polarity), where the Aspect term is the target or entity being discussed, the Opinion term is the sentiment or opinion expressed about the aspect, and Polarity indicates whether the opinion is positive, negative, or neutral. Figure 1 illustrates the ASTE task.
@@ 2.1 ASTE Paradigms
Peng et al. (2020) proposed a pipeline method that divides ASTE tasks into two stages: initially extracting (Aspect, Opinion) pairs and subsequently predicting sentiment polarity. However, pipeline methods typically suffer from error propagation issues Xu et al. (2020). Recent pipeline methods treat ASTE as a Machine Reading Comprehension problem, and develops seq2seq methods such as machine reading comprehension Zhai et al. (2022); Mao et al. (2021); Zou et al. (2024); Chen et al. (2021b). Joint Tagging strategies are remarked by certain Unified Tagging Scheme designs, where elements of a triplet can be extracted simultaneously. ET Xu et al. (2020) introduced a position-awared tagging scheme with a conditional random field module, effectively addressing span overlapping issues. Recent joint paradigm methods have refined the ASTE task with the development of proficient Grid Tagging Schemes (GTS).
@@ 3.1 Overall Framework
An overall description of the training process can be found in Figure 2. Basically, our design can be broken down into the Minimalist Grid Tagging Scheme (GTS) and the Token-level Contrastive Learning Strategy.
@@
Then, the inference phase involves with forming the Minimalist GTS and predicting the corresponding class for each cell. Once the GTS is predicted, it can be decoded by the GTS decoder into triplets in natural language form. The training phase also introduces a novel contrastive learning strategy, in which similar and dissimilar pairs of contextual representations are distinguished. The contrastive loss is then weighted and summed with the tagging loss, which is the classification loss between the predicted and ground-truth tagging schemes.
Our research benefits from the following two closely intertwined aspects: 1) The use of the Minimalist GTS simplifies the learning process by reducing the number of labels, facilitating faster convergence, and seamlessly gearing the contrastive learning. 2) The token-level contrastive learning enhances the model’s ability to distinguish between related and unrelated elements within the input sequence, thereby improving the overall accuracy of the tagging system. For a more detailed description of our algorithm pipeline, see the pseudo-code in Appendix A.1.
@@ 3.2.1 Tagging Scheme Design
As defined by Section 3.1, once an input sentence is encoded into a sequence of contextual representations $\textit{{h}}=\{h_{1},h_{2},...,h_{|\textit{{h}}|}\}$, we form a $|\textit{{h}}|\times|\textit{{h}}|$ matrix, that is, our tagging scheme $\textit{{tag}}_{|\textit{{h}}|\times|\textit{{h}}|}$. As shown in Figure 3, on the rows we mark Aspect tokens by yellow and the columns we mark Opinion tokens by green (positive) and blue (negative). Then the each intersection of these marked rows and columns can uniquely represent an identical sentiment triplet. Thus, each such triplet can be denoted by a 2-D area (submatirx) in the matrix, where Sentiment Polarity is indicated with POS. (positive), NEU. (neutral), or NEG. (negative) in the top-left corner cell of the area, while CTD. indicates the continuation of the pairing relationship within the same region. MSK. (mask) on the diagonal represents masked cells that are not involved in the computation. In Figure 3, an example sentence is tokenized and tagged.
@@ 3.2.2 Tagging Loss
We adopt a tagging loss to guide neural network learning. We concatenate the representation with its transposed form to construct a matrix. Then, we apply the classification head $\mathrm{Cls}$ to the embeddings, followed by the softmax function, to obtain the predicted classification probabilities for each cell:
@@
3.3 Contrstive Learning Strategy
@@ 4.4.1 Comparison to Existing Methods
We evaluate ASTE performance using the widely accepted (Precision, Recall, F1) metrics. The results of the data set $\mathcal{D}_{2}$ are shown in Table 1, while the results of $\mathcal{D}_{1}$ are presented in Appendix Table 8. The best results are highlighted in bold, and the second-best results are underscored. Our proposed method consistently achieves state-of-the-art performance or ranks second in most evaluated cases.
In particular, on the data set $\mathcal{D}_{1}$, the proposed method achieves a substantial improvement of 3. 08% in the F1 score in the 14Lap subset. This improvement is particularly significant given that the highest score on this subset is the lowest among all datasets, showcasing our method’s effectiveness in handling challenging instances. Moreover, on the 14Res subset, our F1 score exceeds 76.00, which, to the best of our knowledge, is the highest reported performance. For the dataset $\mathcal{D}_{2}$, our method outperforms all state-of-the-art approaches by more than 1 percentage point on the 14Res, 14Lap, and 16Res subsets. Only on the 16Res subset does the BDTF method (Zhang et al., 2022b) achieve a slightly better performance.
@@ 5.3 Efficiency Comparison
Table 3 provides another comparative analysis of the tagging schemes. Our method has a compact design with the fewest classes of labels. Moreover, compared with other SOTA baselines, our method does not rely on any additional linguistic information enhancement.
@@ 5.4 Case Study
It is quite interesting to investigate the GPT error cases. The findings reveal that the performance of the GPT model is mixed - while it is able to identify more aspect-opinion (A-O) pairs than the ground truth annotations in some cases, this comes at the cost of a decreased precision. This suggests that the GPT model may be “over-interpreting” the input, making inferences that go beyond what is strictly supported by the text. Furthermore, the GPT model appears to be overly sensitive to the presence of adverbs (such as “very”, “a bit”, etc.) in the input. This sensitivity manifests itself in the model frequently adding or removing adverbs when extracting the opinion components, which further contributes to a decrease in the overall accuracy of the ABSA task.
@@ 6 Conclusion
In this work, we have introduced an elegant and efficient framework for ASTE to achieve SOTA performance. Our approach is built upon two effective components: a new tagging scheme and a novel token-level contrastive learning implementation. The ablation study demonstrates the synergy between these components, reducing the need for complex model designs and external information enhancements.
@@ 7 Limitations
Our method is based on a 2D matrix tagging scheme, where the time complexity for decoding, given the input corpus length $N$, is $O(N^{2})$. This may be unacceptable when $N$ is too large. Additionally, although we have demonstrated our method on commonly used classic English datasets, it should be tested on more natural corpora and for its cross-language capability.
@@ Appendix A Appendix
3:  Similarity matrix:
$\mathbf{Sim}_{|\mathcal{H}|\times|\mathcal{H}|}=-(\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}-\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}^{T})\circ(\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}-\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}^{T})$
where
@@ A.3 Rethinking the GTS
Lemma 2. In a 2D-tagging for ASTE, at least three basic goals must be met: 1) correctly identifying the (Aspect, Opinion) pairs, 2) correctly classifying the sentiment polarity of the pair based on the context, and 3) avoiding boundary errors, such as overlapping [^1] It occurs when one single word belongs to multiple classes in different triplets. , confusion [^2] It occurs when there is a lack of location restrictions so that multiple neighbored candidates can not be uniquely distinguished. , and conflict [^3] It occurs when one single word is composed of multiple tokens, and the predict gives predictions that are not aligned with the word span. . See the proof in Appendix Proof 2.
@@
Figure 6 shows a comprehensive case of our tagging scheme, in which the left matrix is an appearance of our tagging scheme and can be decomposed into two separate components. The middle matrix is the first component, which takes only one tag to locate the up-left beginning of an area, and the second component simply predicts a binary classification to figure out the full area.
@@ A.5 Proof 2:
    Assuming Incorrect Identification: Assume that some (Aspect, Opinion) pairs are incorrectly identified in $M$. This would mean that there exists at least one pair $(i,j)$ where $M[i][j]$ does not represent the actual relationship (Aspect, Opinion) in $S$. This misrepresentation leads to incorrect sentiment analysis results, which contradicts the requirement of the task to provide accurate sentiment analysis, thereby proving that our identification must be correct.
@@ Conclusion
The contraposition approach solidifies that the ASTE tagging strategy in a 2D labeling framework successfully achieves the correct identification of pairs, accurate classification of sentiment, and effective management of boundary errors, as any failure in these aspects leads to contradictions with the task requirements.
@@ A.6 Baselines
few-shots Building upon the zero-shot method, a small number of examples from the training set are added to the prompted sentence: "Perform aspect-based sentiment analysis on the provided text and return triplets as [Aspect, Opinion, Sentiment]. For example: input: {train sentence} output: {train triplets}, … (some other examples). You only need to provide the triplets, no additional explanations are required. The provided text: {sentence}"
@@ A.9 Details Settings of GPT Experients
GPT-3.5 zero-shot GPT-3.5 few-shots GPT-4 zero-shot GPT-4 few-shots
@@
few-shots Suppose you are an expert of aspect-based sentiment analysis. Perform aspect-based sentiment analysis on the provided text and return triplets as [Aspect, Opinion, Sentiment]. For example:
@@ 3. Summarize:
CoT+few-shots Suppose you are an expert of aspect-based sentiment analysis. Please analyze the given text for aspect-based sentiment analysis using the following steps:

### v2

@@ Abstract
Aspect Sentiment Triplet Extraction (ASTE) aims to co-extract the sentiment triplets in a given corpus. Existing approaches within the pretraining-finetuning paradigm tend to either meticulously craft complex tagging schemes and classification heads, or incorporate external semantic augmentation to enhance performance. In this study, we, for the first time, re-evaluate the redundancy in tagging schemes and the internal enhancement in pretrained representations. We propose a method to improve and utilize pretrained representations by integrating a minimalist tagging scheme and a novel token-level contrastive learning strategy. The proposed approach demonstrates comparable or superior performance compared to state-of-the-art techniques while featuring a more compact design and reduced computational overhead. Additionally, we are the first to formally evaluate GPT-4’s performance in few-shot learning and Chain-of-Thought scenarios for this task. The results demonstrate that the pretraining-finetuning paradigm remains highly effective even in the era of large language models. The codebase is available at https://github.com/qiaosun22/MiniConGTS.
11footnotetext: Corresponding Authors
@@ 1 Introduction
Aspect-Based Sentiment Analysis (ABSA) aims to jointly extract opinion terms, aspect terms (targets of the corresponding opinions), and their specific sentiment polarities in a given corpus. In the milestone research by Peng et al. (2020), the compound ABSA subtasks were consolidated into the Aspect Sentiment Triplet Extraction (ASTE) task framework. For each input corpus, ASTE outputs triplets in the form (Aspect, Opinion, Polarity), where the Aspect term is the target or entity being discussed, the Opinion term is the sentiment or opinion expressed about the aspect, and Polarity indicates whether the opinion is positive, negative, or neutral. Figure 1 illustrates the ASTE task.
@@ 2.1 ASTE Paradigms
Peng et al. (2020) proposed a pipeline method that divides ASTE tasks into two stages: initially extracting (Aspect, Opinion) pairs and subsequently predicting sentiment polarity. However, pipeline methods typically suffer from error propagation issues Xu et al. (2020). Recent pipeline methods treat ASTE as a Machine Reading Comprehension problem, and develops seq2seq methods such as machine reading comprehension Zhai et al. (2022); Mao et al. (2021); Zou et al. (2024); Chen et al. (2021b). Joint Tagging strategies are remarked by certain Unified Tagging Scheme designs, where elements of a triplet can be extracted simultaneously. ET Xu et al. (2020) introduced a position-aware tagging scheme with a conditional random field module, effectively addressing span overlapping issues. Recent joint paradigm methods have refined the ASTE task with the development of proficient Grid Tagging Schemes (GTS).
@@ 3.1 Overall Framework
An overall description of the training process can be found in Figure 2. Basically our design can be break down into the Minimalist Grid Tagging Scheme (GTS) and the Token-level Contrastive Learning Strategy.
@@
Then, the inference phase involves with forming the Minimalist GTS and predicting the corresponding class for each cell. Once the GTS is predicted, it can be decoded by the GTS decoder into the triplets in natural language form. The training phase additionally introduces a novel contrastive learning strategy, where similar and dissimilar pairs of contextual representations are distinguished. The contrastive loss is then weighted and summed with the tagging loss, which is the classification loss between the predicted and ground truth tagging schemes.
Our research benefits from the following two closely intertwined aspects: 1) The use of the Minimalist GTS simplifies the learning process by reducing the number of labels, facilitating faster convergence and seamlessly gearing the contrastive learning. 2) The token-level contrastive learning enhances the model’s ability to distinguish between related and unrelated elements within the input sequence, thereby improving the overall accuracy of the tagging system. For a more detailed description for our algorithm pipeline, see the pseudo code in Appendix A.1.
@@ 3.2.1 Tagging Scheme Design
As defined by Section 3.1, once an input sentence is encoded into a sequence of contextual representations $\textit{{h}}=\{h_{1},h_{2},...,h_{|\textit{{h}}|}\}$, we form a $|\textit{{h}}|\times|\textit{{h}}|$ matrix, that is, our tagging scheme $\textit{{tag}}_{|\textit{{h}}|\times|\textit{{h}}|}$. As shown in Figure 3, on the rows we mark Aspect tokens by yellow and the columns we mark Opinion tokens by green (positive) and blue (negative). Then the each intersection of these marked rows and columns can uniquely represent an identical sentiment triplet. Thus, each such triplet can be noted by a 2-D area (submatirx) in the matrix, where Sentiment Polarity is indicated with POS. (positive), NEU. (neutral), or NEG. (negative) in the top-left corner cell of the area, while CTD. indicates the continuation of the pairing relationship within the same region. MSK. (mask) on the diagonal represents masked cells that are not involved in the computation. In Figure 3, an example sentence is tokenized and tagged.
@@ 3.2.2 Tagging Loss
We adopt a tagging loss to guide neural network learning.
We concatenate the representation with its transposed form to construct a matrix. Then, we apply the classification head $\mathrm{Cls}$ to the embeddings, followed by the softmax function, to obtain the predicted classification probabilities for each cell:
@@
3.3 Contrastive Learning Strategy
@@ 4.4.1 Comparison to Existing Methods
We evaluate ASTE performance using the widely accepted (Precision, Recall, F1) metrics. Results of the dataset $\mathcal{D}_{2}$ are shown in Table 1, while the results of $\mathcal{D}_{1}$ are presented in Appendix Table 8. The best results are highlighted in bold, and the second-best results are underscored. Our proposed method consistently achieves state-of-the-art performance or ranks second in most evaluated cases.
Notably, on dataset $\mathcal{D}_{1}$, the proposed method achieves a substantial 3.08% improvement in F1 score on the 14Lap subset. This improvement is particularly significant given that the highest score on this subset is the lowest among all datasets, showcasing our method’s effectiveness in handling challenging instances. Moreover, on the 14Res subset, our F1 score exceeds 76.00, which, to the best of our knowledge, is the highest reported performance. For dataset $\mathcal{D}_{2}$, our method outperforms all state-of-the-art approaches by over 1 percentage point on the 14Res, 14Lap, and 16Res subsets. Only on the 16Res subset does the BDTF method (Zhang et al., 2022b) achieve a slightly better performance.
@@ 5.3 Efficiency Comparison
Table 3 provides another comparative analysis of tagging schemes. Our method has a compact design with the fewest classes of labels. What’s more, compared with other SOTA baselines, our method does not rely on any additional linguistic information enhancement.
@@ 5.4 Case Study
It is quite interesting to investigate the GPT’s error cases. The findings reveal that the performance of the GPT model is mixed - while it is able to identify more aspect-opinion (A-O) pairs than the ground truth annotations in some cases, this comes at the cost of a decreased precision. This suggests that the GPT model may be “over-interpreting” the input, making inferences that go beyond what is strictly supported by the text. Furthermore, the GPT model appears to be overly sensitive to the presence of adverbs (such as “very”, “a bit”, etc.) in the input. This sensitivity manifests in the model frequently adding or removing adverbs when extracting the Opinion components, which further contributes to a decrease in the overall accuracy of the ABSA task.
@@ 6 Conclusion
In this work, we have introduced an elegant and efficient framework for ASTE, achieving SOTA performance. Our approach is built upon two effective components: a new tagging scheme and a novel token-level contrastive learning implementation. The ablation study demonstrates the synergy between these components, reducing the need for complex model designs and external information enhancements.
@@ 7 Limitations
Our method is based on a 2D-matrix tagging scheme, where the time complexity for decoding, given the input corpus length $N$, is $O(N^{2})$. This may be unacceptable when $N$ is too large. Additionally, although we have demonstrated our method on commonly used classic English datasets, it should be tested on more natural corpora and for its cross-language capability.
@@ 8 Ethics & Potential Risks Statement
9 Acknowledgement
This work was partially supported by the National Key R&D Program of China under Grant No. 2022ZD0160101. The authors would also like to thank Shanghai Artificial Intelligence Laboratory for their support and resources, as well as colleagues from Shanghai Jiao Tong University for valuable discussions. This work was conducted during the internships of Qiao Sun, Liujia Yang, and Minghao Ma at the Shanghai Artificial Intelligence Laboratory.
@@ Appendix A Appendix
3:  Similarity matrix: $\mathbf{Sim}_{|\mathcal{H}|\times|\mathcal{H}|}=$ (||×||||×$-(\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}-\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}^{T})\circ(\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}-\mathcal{H}_{|\mathcal{H}|\times|\mathcal{H}|}^{T})$ where
@@ A.3 Rethinking the GTS
Lemma 2. In a 2D-tagging for ASTE, at least three basic goals must be met: 1) correctly identifying the (Aspect, Opinion) pairs, 2) correctly classifying the sentiment polarity of the pair based on the context, and 3) avoiding boundary errors, such as overlapping***It occurs when one single word belongs to multiple classes in different triplets. , confusionIt occurs when there is a lack of location restrictions so that multiple neighbored candidates can not be uniquely distinguished. , and conflictIt occurs when one single word is composed of multiple tokens, and the predict gives predictions that are not aligned with the word span. . See the proof in Appendix Proof 2.
@@
Figure 6 shows a comprehensive case of our tagging scheme, in which the left matrix is an appearance of our tagging scheme, and it can be decomposed into two separate components. The middle matrix is the first component, which takes only one tag to locate the up-left beginning of an area, and the second component simply predicts a binary classification to figure out the full area.
@@ A.5 Proof 2:
    Assuming Incorrect Identification: Assume that some (Aspect, Opinion) pairs are incorrectly identified in $M$. This would mean that there exists at least one pair $(i,j)$ where $M[i][j]$ does not represent the actual (Aspect, Opinion) relationship in $S$. This misrepresentation leads to incorrect sentiment analysis results, which contradicts the requirement of the task to provide accurate sentiment analysis, thereby proving that our identification must be correct.
@@ Conclusion
The contraposition approach solidifies that the tagging strategy for ASTE in a 2D labeling framework successfully achieves the correct identification of pairs, accurate classification of sentiment, and effective management of boundary errors, as any failure in these aspects leads to contradictions with the task requirements.
@@ A.6 Baselines
few-shot Building upon the zero-shot method, a small number of examples from the training set are added to the prompted sentence: "Perform aspect-based sentiment analysis on the provided text and return triplets as [Aspect, Opinion, Sentiment]. For example: input: {train sentence} output: {train triplets}, … (some other examples). You only need to provide the triplets, no additional explanations are required. The provided text: {sentence}". We utilized 5-shot, 10-shot, and 20-shot methods, all randomly sampled from the training set. The results indicate that the 5-shot method performed the best, while the performances of the 10-shot and 20-shot methods showed a decline. The tables presents the output results for the 5-shot method.
@@ A.9 Details Settings of GPT Experients
GPT-3.5 zero-shot GPT-3.5 few-shot GPT-4 zero-shot GPT-4 few-shot
@@
few-shot Suppose you are an expert of aspect-based sentiment analysis. Perform aspect-based sentiment analysis on the provided text and return triplets as [Aspect, Opinion, Sentiment]. For example:
@@ 3. Summarize:
CoT+few-shot Suppose you are an expert of aspect-based sentiment analysis. Please analyze the given text for aspect-based sentiment analysis using the following steps: